# Classification of thin Jordan schemes

### v1

@@ 4.2. Thin regular Jordan schemes are autonomous
In this subsection we will prove the statement in the title. To formulate it exactly we change slightly some notation for this subsection. Namely, if $(S,\diamond)$ is a loop, then we’ll write here $\hat{s}$ instead of $L_{s}$ used previously. The set of all permutations $\{\hat{s}\}_{s\in S}$ will be denoted as $\hat{S}$.
Now we are ready to formulate the main result of this subsection.
@@
Assume towards a contradiction that $\mathfrak{S}$ is an algebraic fusion of a coherent configuration $\mathfrak{C}=(S,C)$ w.r.t. to a certain subgroup $\Phi\leq\mathsf{TAut}(\mathfrak{C})$. Then each $\hat{s}=\Phi(c)$ for each $s\in S$ and every $c\in C$ contained in $\hat{s}$. In particular, this implies that each $c\in C$ is thin. By [5, Exercise 2.7.35] every such configuration is Schurian, i.e. there exists $H\leq{\sf Sym}(S)$ such that $C$ is a $2$-orbit set of $H$. Since each $2$-orbit of $H$ is thin, $H$ is semiregular. By Proposition 4.3 $\mathsf{JAut}(\mathfrak{C})=\mathsf{TAut}(\mathfrak{C})$. Hence the subgroup $\Phi_{0}:=\Phi\cap\mathsf{AAut}(\mathfrak{C})$ has index at most two in $\Phi$. Since $\Phi_{0}\leq\mathsf{AAut}(\mathfrak{C})$, the rainbow $^{\Phi_{0}}\mathfrak{C}=(S,C_{0})$ with $C_{0}:=\{\Phi_{0}(c)\}_{c\in C}$ is a coherent configuration. Thus $\hat{S}={}^{\Phi}C\sqsubseteq C_{0}\sqsubseteq C$. Since $\mathfrak{S}$ is not a co.co., the inclusion $\hat{S}\sqsubseteq C_{0}$ is proper and $\Phi:\Phi_{0}]=2$. In particular, $\Phi_{0}$ is normal in $\Phi$, which, in turn, implies that each $\Phi$-orbit on $C$ is either a $\Phi_{0}$-orbit or a disjoint union of two $\Phi_{0}$-orbits of the same size. Therefore, each $\hat{s}\in\hat{S}$ is either a basic relation of $^{\Phi_{0}}\mathfrak{C}$ or a union of two basic relations of $^{\Phi_{0}}\mathfrak{C}$ of the same size. Since $^{\Phi_{0}}\mathfrak{C}$ and $^{\Phi}\mathfrak{C}$ are distinct, there exists $\hat{s}\in\hat{S}$ which splits into a disjoint union of two basic relations of $^{\Phi_{0}}\mathfrak{C}$ of equal size, say $\hat{s}=c_{1}\cup c_{2}$. It follows from $|c_{1}|=|c_{2}|=|S|/2$ that both $c_{1}$ and $c_{2}$ are non-regular basic relations of $^{\Phi_{0}}\mathfrak{C}$. Therefore, $^{\Phi_{0}}\mathfrak{C}$ is a non-homogeneous co.co., that is, each of its basic relations is non-regular. This implies that all b.relations of $^{\Phi_{0}}\mathfrak{C}$ have size $|S|/2$.
@@
Notice, that the coherent closure $\mathfrak{T}=(S,T)$ of $\mathfrak{S}$ satisfies $S\sqsubset T\sqsubseteq C_{0}$. By Proposition 4.2 the basic relations of $T$ coincide with $2$-orbits of the right nucleous $N_{\rho}$ of the loop $(S,\diamond)$. In particular, $|t|=|N_{\rho}|$ for each $t\in T$. It follows from Proposition 3.16 that $N_{\rho}=Z(S)$, and, therefore, $|N_{\rho}|=|S|/8$. Since each $t\in T$ is a union of basic relations of $C_{0}$, a basic relation $c\in C_{0}$ has cardinality at most $|S|/8$, contrary to $|c|=|S|/2$. A contradiction. ∎
5. Appendix
Proposition 5.1.
Let $\mathcal{J}\subseteq M_{\Omega}({\mathbb{F}})$ be a Jordan subalgebra, i.e. $\mathcal{J}\star\mathcal{J}\subseteq\mathcal{J}$. Then for any three elements $a,b,c\in\mathcal{J}$ it holds that $acb+bca\in\mathcal{J}$. In particular, $aca\in\mathcal{J}$
Proof.
The inclusion $acb+bca\in\mathcal{J}$ follows from the identity $a\star(b\star c)+b\star(a\star c)-(a\star b)\star c=\tfrac{1}{2}(acb+bca)$. ∎
We do not know whether the above statement can be generalized for $4$ and more elements. More precisely, is it true that for any $n$ elements $a_{1},...,a_{n}\in\mathcal{J}$ the sum $a_{1}\cdots a_{n}+a_{n}\cdots a_{1}$ also belongs to $\mathcal{J}$?
Meanwhile, we have a much weaker statement.
Proposition 5.2.
Let $\mathcal{J}\subseteq M_{\Omega}({\mathbb{F}})$ be a $\star$-subalgebra. Then for any $n$ elements $a_{1},...,a_{n}\in\mathcal{J}$ their symmetrization $S(a_{1},...,a_{n}):=\sum_{\pi\in S_{n}}a_{\pi(1)}\cdots a_{\pi(n)}$ belongs to $\mathcal{J}$
Proof.
Induction on $n$. We know that the claim holds for $n=2$. Suppose that it holds for $n\geq 2$. Let $a_{1},\dots,a_{n+1}\in\mathcal{J}$. Then
$2S(a_{1},\dots,a_{n+1})$ $=\sum_{i=1}^{n+1}a_{i}S(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n+1})+\sum_{i=1}^{n+1}S(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n+1})a_{i}$
$=2\sum_{i=1}^{n+1}a_{i}\star S(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n+1}).$
By induction hypothesis, for each $i\in\{1,\dots,n+1\}$ we have $S(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n+1})\in\mathcal{J}$. Thus, $S(a_{1},\dots,a_{n+1})\in\mathcal{J}$. ∎

### v2

@@ 4.2. Thin regular Jordan schemes are autonomous
Before proceeding to fulfill the task reflected in the subsection title, we make some small notational changes. Namely, elements of a loop $(S,\diamond)$ shall now be denoted by $\hat{s}$ rather than $\ell_{s}$, which was our previously chosen notation. Also, the set of all permutations $\{\hat{s}\}_{s\in S}$ will be denoted by $\hat{S}$.
We are now equipped to prove the main result of this subsection.
@@
By way of contradiction, assume $\mathfrak{S}$ is the algebraic fusion of a coherent configuration $\mathfrak{C}=(S,C)$ with respect to a certain subgroup $\Phi\leq\mathsf{JAut}(\mathfrak{C})$. Then $\hat{s}=\Phi(c)$ for every $s\in S$ and $c\in C$ contained in $\hat{s}$. In particular, this implies that each $c\in C$ is thin. By [5, Exercise 2.7.35], every such configuration is Schurian, i.e. there exists $H\leq{\sf Sym}(S)$ such that $C$ is the set of $2$-orbits of $H$. Since every $2$-orbit of $H$ is thin, we have that $H$ is semiregular. Thus by Proposition 4.3, $\mathsf{JAut}(\mathfrak{C})=\mathsf{TAut}(\mathfrak{C})$. Hence the subgroup $\Phi_{0}:=\Phi\cap\mathsf{AAut}(\mathfrak{C})$ has index at most two in $\Phi$. Moreover, since $\Phi_{0}\leq\mathsf{AAut}(\mathfrak{C})$, the rainbow $^{\Phi_{0}}\mathfrak{C}=(S,C_{0})$ where $C_{0}:=\{\Phi_{0}(c)\}_{c\in C}$ is a coherent configuration. Thus $\hat{S}={}^{\Phi}C\sqsubseteq C_{0}\sqsubseteq C$. Since $\mathfrak{S}$ is not a coherent configuration, the inclusion $\hat{S}\sqsubseteq C_{0}$ must be proper, and $\Phi:\Phi_{0}]=2$. In particular, $\Phi_{0}$ is normal in $\Phi$, which in turn implies that each $\Phi$-orbit on $C$ is either a $\Phi_{0}$-orbit or a disjoint union of two $\Phi_{0}$-orbits of equal size. Therefore, each $\hat{s}\in\hat{S}$ is either a basic relation of $^{\Phi_{0}}\mathfrak{C}$ or the union of two basic relations of $^{\Phi_{0}}\mathfrak{C}$ of equal size. Since $^{\Phi_{0}}\mathfrak{C}$ and $^{\Phi}\mathfrak{C}$ are distinct, there exists $\hat{s}\in\hat{S}$ which splits into a disjoint union of two basic relations of $^{\Phi_{0}}\mathfrak{C}$ of equal size, say $\hat{s}=c_{1}\cup c_{2}$. It follows from $|c_{1}|=|c_{2}|=|S|/2$ that both $c_{1}$ and $c_{2}$ are nonregular basic relations of $^{\Phi_{0}}\mathfrak{C}$. Therefore $^{\Phi_{0}}\mathfrak{C}$ is a nonhomogeneous coherent configuration, that is, each of its basic relations is nonregular. This in turn implies that all basic relations of $^{\Phi_{0}}\mathfrak{C}$ have size $|S|/2$.
@@
Observe that the coherent closure $\mathfrak{T}=(S,T)$ of $\mathfrak{S}$ satisfies $\hat{S}\sqsubset T\sqsubseteq C_{0}$. By Proposition 4.2, the basic relations of $T$ coincide with the $2$-orbits of the right nucleus $N_{\rho}$ of the loop $(S,\diamond)$. In particular, $|t|=|N_{\rho}|$ for each $t\in T$. It follows from Proposition 3.16 that $N_{\rho}=Z(S)$, and therefore $|N_{\rho}|=|S|/8$. Since each $t\in T$ is the union of basic relations of $C_{0}$, a basic relation $c\in C_{0}$ has cardinality at most $|S|/8$, contrary to $|c|=|S|/2$. As this is an obvious contradiction, the proof is complete. ∎
Acknowledgment
The authors wish to thank Misha Klin for helpful discussions and suggestions.